OpenStudy (anonymous):

How many grams of magnesium metal will react completely with 6.3 liters of 4.5 M HCl? Show all of the work needed to solve this problem. Mg (s) + 2 HCl (aq) yields MgCl2 (aq) + H2 (g)

OpenStudy (abmon98):

\[Mg(s)+2HCL(aq)\rightarrow MgCl2+H2(g) \] Number of moles=Volume(l)*Concentration(mol/l) Number of moles of HCL=6.3*4.5=28.35 moles of HCl Look at the mole ratio according to the balanced chemical equation. 1 Mg mole reacts with 2 mole of HCl to produce 1 mole of both H2 and MgCl2 1:2:1:1 now lets in account the mole ratio between the two reactants only. 1:2 x:28.35 x=14.175 Use Number of moles=Mass(g)/Molar Mass(g/mol) Atomic weight of Mg:24

OpenStudy (anonymous):

Mg + 2 HCl → H2 + MgCl2 (6.3 L) x (4.5 mol/L HCl) x (1 mol Mg / 2 mol HCl) x (24.30506 g Mg/mol) = 345 g Mg